# Phase Selection and Analysis for Multi-frequency Multi-user RIS Systems Employing Subsurfaces in Correlated Ricean and Rayleigh Environments

### v1

@@ III Phase Selection Methods
Partitioning the $N$ RIS elements into $K$ subsurfaces and independently designing each subsurface for one user allows for low complexity and effective RIS phase selection methods, which is especially important for the high frequency environments it is envisaged RIS will be used in. This section outlines three methods that utilise these principles: a simple approach that uses the SU optimal phase selection method for when $\mathbf{H}_{\mathrm{rb}}$ is rank-1, a sub-optimal extension of this for when $\mathbf{H}_{\mathrm{rb}}$ has a scattered component, and a higher-performing but more complex iterative version.
@@ III-A LoS Phase Selection Method
As we first proposed in [12], the phases of a subsurface designed for the $k$-th user can be set according to the SU method set out in [13], so that
@@
The phases set for each user are optimal for that user when $\mathbf{H}_{\mathrm{rb}}^{(k)}$ is purely LoS, and close to optimal when $\mathbf{H}_{\mathrm{rb}}^{(k)}$ has a strong LoS component. This method leads to a $K$-fold reduction in the required CSI compared to typical RIS designs, due to only requiring the estimation of $N_{k}$ rather than $N$ channel elements per user. There are no complex matrix operations required and a simple MF receiver can be used. These advantages make this approach a very attractive option for implementation.
@@ III-B Extended Subsurface Design for NLoS Systems
The design in (12) and (13) works well for strongly LoS channels. However, if there is a strong scattered component, a different approach provides better results. As there is no optimal closed form SU phase selection method when $\mathbf{H}^{(k)}_{\mathrm{rb}}$ is not rank 1, we propose the following adaptation to the method outlined in Sec. III-A. As we proposed in [12], we take the singular value decomposition of $\mathbf{H}^{(k)}_{\mathrm{rb},k}$, so that $\mathbf{H}^{(k)}_{\mathrm{rb},k}=\mathbf{UDV^{\dagger}}$, where $\mathbf{D}=\mathrm{diag}(D_{11},D_{22},\dots,D_{mm})$, $m=\min(M,N_{k})$ and the singular values are $D_{11}\geq D_{22}\geq...\geq D_{mm}$. The strongest rank 1 component of $\mathbf{H}^{(k)}_{\mathrm{rb},k}$ is $D_{11}\mathbf{u}_{k}\mathbf{v}_{k}^{\dagger}$, where $\mathbf{u}_{k}$ and $\mathbf{v}_{k}$ are the leading left and right singular vectors, respectively. This has the same structure as a pure LoS channel, so the same design can be used by considering just the strongest component, giving,
@@ III-C Iterative Phase Alignment
The subsurface methods outlined in Secs. III-A and III-B select phases to maximise the norm of the subsurface’s RIS link, $\mathbf{H}^{(k)}_{\mathrm{rb},k}\mathbf{\Phi}_{k}\mathbf{h}^{(k)}_{\mathrm{ur},k}$, and then rotate these phases to align with the non-RIS-controlled direct link, $\mathbf{h}^{(k)}_{\mathrm{d}}$. It is assumed that each subsurface is designed independently, with no knowledge of any phases that have been set for other users. If we assume that the subsurfaces are set sequentially, we can improve these methods by treating previously set subsurfaces as fixed channels. After selecting phases to maximise the RIS link, they can be rotated to align with a combination of the direct path and the RIS paths that have already been set for other UEs. This is done by including this information into $\nu_{k}$ (LoS case) or $\omega_{k}$ (NLoS case). The subsurfaces are set from the weakest UE-RIS channel to the strongest, so that the most impactful channels can benefit from the most information. This process is detailed in Algorithm 1, where index()($\cdot$)( ⋅ ) refers to a function that returns the index of elements in a list and sort()($\cdot$)( ⋅ ) refers to a function that reorders elements of a list from smallest to largest.
This iterative method improves mean SNR as shown in Sec. V-A at the cost of a significant increase in CSI requirements at the BS. $\mathbf{h}_{\mathrm{ur}}^{(k)}$ contains $N$ channel elements, with $\frac{N}{K}$ corresponding to those impacted by a specific subsurface. The iterative method requires the computation of the magnitude of a user’s UE-RIS link to rank channels based on their impact, so knowledge of the amplitude of all $N$ channel elements is required. In contrast, the methods in Secs. III-A and III-B only consider the impact on a user from each subsurface, so only the subsurface specific $\frac{N}{K}$ elements are required. The benefits of the iterative method vary depending on channel conditions such as correlation and link power, as will be shown in Section V-A.
@@
[…]
     for $l=1$ to $k-1$ do
[…]
@@
[…]
@@ IV Analysis
In this section, we focus on an analysis of the mean SNR as prior work [13, 12, 18] has shown it to be tractable, to provide tight rate bounds via $\log_{2}(1+\mathbb{E}[\mathrm{SNR}])$ and to deliver system insights. In particular, we derive the mean SNR for a correlated Ricean environment, where $\mathbf{h}_{\mathrm{d}}$, $\mathbf{H}_{\mathrm{rb}}$ and $\mathbf{h}_{\mathrm{ur}}$ all consist of a rank-1 LoS and a scattered component. We consider the important case where $\mathbf{H}_{\mathrm{rb}}$ has a strong, but not pure, LoS component. This leads to a stable, slow moving RIS-BS channel, allowing the SD to be used and achieve close to optimal performance for a given subsurface [19]. Therefore, $\mathbf{\Phi}$ in (12) is used in this section. This general case will then be simplified to two special cases where $\mathbf{h}_{\mathrm{d}}$ and $\mathbf{h}_{\mathrm{ur}}$ are correlated Rayleigh channels; in the first $\mathbf{H}_{\mathrm{rb}}$ is correlated Ricean, and in the second it is a rank-1 LoS channel.
The received signal $\mathbf{r}_{k}$ from (2) can be rewritten as
$\mathbf{r}_{k}=(\mathbf{h}^{(k)}_{\mathrm{d}}+\mathbf{f}_{k}+\mathbf{g}_{k})s_{k}+\mathbf{n}_{k},$ (16)
@@ IV-A Mean SNR for Correlated Ricean Systems
$\mathbb{E}[\mathbf{h}^{(k)\dagger}_{\mathrm{d}}\mathbf{f}_{\mathrm{k}}]=\frac{N_{k}\pi||\mathbf{R}^{(k)1/2}_{\mathrm{d}}\mathbf{a}^{(k)}_{\mathrm{b}}||\sqrt{\!\beta^{(k)}_{\mathrm{d}}\!\beta^{(k)}_{\mathrm{rb}}\!\beta^{(k)}_{\mathrm{ur}}}}{4}\\ \times\eta_{\mathrm{rb}}^{(k)}\zeta_{\mathrm{d}}^{(k)}\zeta_{\mathrm{ur}}^{(k)}L_{1/2}\!\left(\!-\kappa_{\mathrm{ru}}^{(k)}\!\right)\!L_{1/2}\!\!\left(\frac{-\kappa_{\mathrm{d}}^{(k)}|\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{a}_{\mathrm{d}}^{(k)}|^{2}}{||\mathbf{R}^{(k)1/2}_{\mathrm{d}}\mathbf{a}_{\mathrm{b}}^{(k)}||^{2}}\!\right).$ (23)
@@
×$\times\!{}_{1}F_{1}\!\left(\!\frac{1}{2},2,-\frac{|\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{a}_{\mathrm{d}}^{(k)}|^{2}\kappa_{\mathrm{d}}^{(k)}}{\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{R}_{\mathrm{d}}^{(k)}\mathbf{a}_{\mathrm{b}}^{(k)}}\!\right)\!\sum_{s\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}\bigg{[}\!\Big{(}\mathbf{a}^{(k)}_{\mathrm{ur},s,j}\!-\!\rho^{(k)}_{ks}$
×$\times\mathbf{a}_{\mathrm{ur},k,i}^{(k)}\Big{)}{}_{1}F_{1}\!\!\left(\!\frac{3}{2},1,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!+\!\frac{3}{2}\rho^{(k)}_{ks}{}_{1}F_{1}\!\!\left(\!\frac{5}{2},2,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}}\!\bigg{]}$
×$\times\!\left(\!\kappa_{\mathrm{rb}}^{(k)}\mathbf{a}^{(k)}_{\mathrm{r},k,i}\mathbf{a}^{(k)*}_{\mathrm{r},s,j}\!+\!r_{ks}^{(k)}\right)\!(\mathbf{C}(s))_{j,j}\,\mathrm{e}^{-j\angle\mathbf{a}_{\mathrm{r},k,i}^{(k)}},$ (30)
@@
×$\times\!\cos(n\phi_{f})\frac{\epsilon_{n}|\rho^{(k)}_{kk}|^{2m-n}}{m!(m-n)!(n!)^{2}}\!\left(\!\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}-2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\!\right)^{\!\!n}\!\!$
×$\times\!\Gamma^{2}\!\!\left(\!\!m\!+\!\frac{3}{2}\!\right)\!{}_{1}F_{1}^{2}\!\!\left(\!m\!+\!\frac{3}{2},n\!+\!1,\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}-2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\right)\!,$ (35)
@@
×$\quad\times\mathrm{e}^{-j\angle\mathbf{a}_{\mathrm{r},s,i}^{(s)}}\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{r},s,j}^{(s)}}\Big{(}\kappa_{\mathrm{ur}}^{(k)}\mathbf{a}_{\mathrm{ur},s,i}^{(k)*}\mathbf{a}_{\mathrm{ur},s,j}^{(k)}+\rho^{(k)}_{ss}\Big{)}\Big{(}\kappa_{\mathrm{rb}}^{(k)}$
×$\quad\times\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},s,j}^{(k)*}+r^{(k)}_{ss}\Big{)}\!+\!\!\sum_{s\neq t\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}(\mathbf{C}^{*}\!(s)\!)_{i,i}(\mathbf{C}(t)\!)_{j,j}$
×$\quad\times\!\Big{(}\!\kappa_{\mathrm{ur}}^{(k)}\mathbf{a}_{\mathrm{ur},s,i}^{(k)*}\mathbf{a}_{\mathrm{ur},t,j}^{(k)}\!+\!\rho^{(k)}_{st}\!\Big{)}\!\Big{(}\!\kappa_{\mathrm{rb}}^{(k)}\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},t,j}^{(k)*}\!+\!r^{(k)}_{st}\!\Big{)}\!\Bigg{]}\!,$ (36)
@@
×$\times\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}\frac{\epsilon_{n}\kappa_{0}^{2m+n}}{2^{m+n}m!\Gamma(m\!+\!n\!+\!1)}\left(\mathcal{F}(3)\mathrm{e}^{jn\phi}\!+\!\mathcal{F}(1)\mathrm{e}^{-jn\phi}\right)\!,$ (37)
@@
×$\!\!\!\times\!\!{}_{1}F_{1}^{2}\!\!\left(\!\!n\!+\!m\!+\!\frac{x}{2},\!n\!-\!1\!+\!x,\!\frac{\zeta^{2}\!\kappa_{\mathrm{ur}}^{(s)}\!\!\left(\!1\!+\!|\rho_{ss}^{(s)}|^{2}\!-\!2\mu_{c}\!\right)\!\!\!\left(\!1\!-\!\rho_{ss}^{(s)2\!}\right)}{4\left(1\!+\!\kappa_{\mathrm{ur}}^{(s)}\right)}\!\!\right)\!\!,$ (38)
@@
Therefore, combining (20), (23), (26), (30), (33) and (36) with (19) gives the complete mean SNR.
@@
$\mathbb{E}[\mathbf{h}^{(k)\dagger}_{\mathrm{d}}\mathbf{f}_{\mathrm{k}}]=\frac{N_{k}\pi||\mathbf{R}^{(k)1/2}_{\mathrm{d}}\mathbf{a}^{(k)}_{\mathrm{b}}||\eta_{\mathrm{rb}}^{(k)}\sqrt{\!\beta^{(k)}_{\mathrm{d}}\!\beta^{(k)}_{\mathrm{rb}}\!\beta^{(k)}_{\mathrm{ur}}}}{4},$ (40)

### v2

@@ III Phase Selection Methods
Partitioning the $N$ RIS elements into $K$ subsurfaces each designed independently for a specific UE enables low-complexity and efficient RIS phase selection. This section outlines three methods that utilise this concept: a simple approach that uses the SU optimal phase selection method for when $\mathbf{H}_{\mathrm{rb}}$ is rank-1, a sub-optimal extension of this for when $\mathbf{H}_{\mathrm{rb}}$ has a scattered component, and a higher-performing but more complex iterative version.
@@ III-A LoS Phase Selection Method
As we first proposed in [16], the phases of a subsurface designed for the $k$-th UE can be set according to the SU method set out in [18], so that
@@
The phases set for each UE are optimal for that UE when $\mathbf{H}_{\mathrm{rb}}^{(k)}$ is purely LoS, and close to optimal when $\mathbf{H}_{\mathrm{rb}}^{(k)}$ has a strong LoS component. This method leads to a $K$-fold reduction in the required CSI compared to typical RIS designs, due to only requiring the estimation of $N_{k}$ rather than $N$ channel elements per user. There are no complex matrix operations required and a simple MF receiver can be used. These advantages make this approach a very attractive option for implementation.
@@ III-B Extended Subsurface Design for NLoS Systems
The design in (12) and (13) works well for strongly LoS channels. However, if there is a strong scattered component, a different approach provides better results. As there is no optimal closed form SU phase selection method when $\mathbf{H}^{(k)}_{\mathrm{rb}}$ is not rank-1, we propose the following adaptation to the method outlined in Sec. III-A, referred to as the extended subsurface design (ESD). As we proposed in [16], we take the singular value decomposition (SVD) of $\mathbf{H}^{(k)}_{\mathrm{rb},k}$, so that $\mathbf{H}^{(k)}_{\mathrm{rb},k}=\mathbf{UDV^{\dagger}}$, where $\mathbf{D}=\mathrm{diag}(D_{11},D_{22},\dots,D_{mm})$, $m=\min(M,N_{k})$ and the singular values are $D_{11}\geq D_{22}\geq...\geq D_{mm}$. The strongest rank-1 component of $\mathbf{H}^{(k)}_{\mathrm{rb},k}$ is $D_{11}\mathbf{u}_{k}\mathbf{v}_{k}^{\dagger}$, where $\mathbf{u}_{k}$ and $\mathbf{v}_{k}$ are the leading left and right singular vectors, respectively. This has the same structure as a pure LoS channel, so the same design can be used by considering just the strongest component, giving,
@@ III-C Iterative Phase Alignment
The subsurface methods outlined in Secs. III-A and III-B assume that each subsurface is designed independently, with no knowledge of phases that have been set for other users. Since the reflections produced by the other subsurfaces manifest as random scattering for a given user, setting the subsurfaces sequentially allows this randomness to be incorporated into the design and used to inform the configuration of the remaining subsurfaces.
The methods in the previous sections select phases to maximize the norm of the subsurface’s RIS link, $\mathbf{H}^{(k)}_{\mathrm{rb},k}\mathbf{\Phi}_{k}\mathbf{h}^{(k)}_{\mathrm{ur},k}$, and then rotate these phases to align with the non-RIS-controlled direct link, $\mathbf{h}^{(k)}_{\mathrm{d}}$. The ISD instead involves selecting phases to maximize the RIS link, and then aligning these with a combination of the direct path and the RIS paths that have already been set for other UEs. This is done by including this information into $\nu_{k}$ (LoS case) or $\omega_{k}$ (NLoS case). The subsurfaces are set from the weakest UE-RIS channel to the strongest, so that the most impactful channels can benefit from the most information. This process is detailed in Algorithm 1, where index()($\cdot$) refers to a function that returns the index of elements in a list and sort()($\cdot$) refers to a function that reorders elements of a list from smallest to largest.
This iterative method improves the mean SNR as shown in Sec. V-A at the cost of a significant increase in CSI requirements at the BS. $\mathbf{h}_{\mathrm{ur}}^{(k)}$ contains $N$ channel elements, with $\frac{N}{K}$ corresponding to those impacted by a specific subsurface. The iterative method requires the computation of the magnitude of a user’s UE-RIS link to rank channels based on their impact, so knowledge of the amplitude of all $N$ channel elements is required. In contrast, the methods in Secs. III-A and III-B only consider the impact on a UE from each subsurface, so only the subsurface specific $\frac{N}{K}$ elements are required. The benefits of the iterative method vary with channel conditions such as correlation and link power, as will be shown in Sec. V-A.
@@
  $\mathit{index\_k}=\mathit{user\_order}(k)$
  $\mathit{fixed}=0$
  for $l=1$ to k1k-1 do
   $\mathit{index\_l}=\mathit{user\_order}(l)$
   $\mathit{fixed}=\mathit{fixed}+\mathbf{H}^{(\mathit{index\_k})}_{\mathrm{rb},\mathit{index\_l}}\mathbf{\Phi}_{\mathit{index\_l}}\mathbf{h}^{(\mathit{index\_k})}_{\mathrm{ur},\mathit{index\_l}}$
@@
  $\mathit{direct}=\mathbf{h}^{(\mathit{index\_k})}_{\mathrm{d}}+\mathit{fixed}$
  $\nu_{\mathit{index\_k}}=\dfrac{\mathbf{a}_{\mathrm{b}}^{(\mathit{index\_k})\dagger}\times\mathit{direct}}{|\mathbf{a}_{\mathrm{b}}^{(\mathit{index\_k})\dagger}\times\mathit{direct}|}$
  $\mathbf{\Phi}_{\mathrm{\mathit{index\_k}}}=\nu_{\mathit{index\_k}}\,\mathrm{diag}\left(\mathrm{e}^{j\left(\angle\mathbf{a}^{(\mathit{index\_k})}_{\mathrm{r}}-\angle\mathbf{h}^{(\mathit{index\_k})}_{\mathrm{ur},\mathit{index\_l}}\right)}\right)$
@@ IV Analysis
In this section, we focus on an analysis of the mean SNR as prior work [18, 16, 24] has shown it to be tractable and to deliver system insights. The accuracy of this rate bound was investigated for the SD in our work in [16], and it was found to be tight for all values. In particular, we derive the mean SNR for a correlated Ricean environment, where $\mathbf{h}_{\mathrm{d}}$, $\mathbf{H}_{\mathrm{rb}}$ and $\mathbf{h}_{\mathrm{ur}}$ all consist of a rank-1 LoS and a scattered component. We consider the important case where $\mathbf{H}_{\mathrm{rb}}$ has a strong, but not pure, LoS component. This leads to a stable, slow-moving RIS-BS channel, allowing the SD to be used with close to optimal performance for a given subsurface [25]. Unfortunately, as far as we know, a closed-form result for the mean SNR is not available using the ESD. However, the SD remains highly accurate in the considered scenario. This enables the derivation of closed-form results, which are particularly valuable for understanding how system behavior arises from underlying mathematical structure. Therefore, $\mathbf{\Phi}$ in (12) is used in this section. This general case is then simplified to two special cases where $\mathbf{h}_{\mathrm{d}}$ and $\mathbf{h}_{\mathrm{ur}}$ are correlated Rayleigh channels; in the first $\mathbf{H}_{\mathrm{rb}}$ is correlated Ricean, and in the second it is a rank-1 LoS channel. The received signal $\mathbf{r}_{k}$ in (2) can be rewritten as
$\mathbf{r}_{k}=\left(\mathbf{h}^{(k)}_{\mathrm{d}}+\mathbf{f}_{k}+\mathbf{g}_{k}\right)s_{k}+\mathbf{n}_{k},$ (16)
@@ IV-A Mean SNR for Correlated Ricean Systems
[…]
Remark 1
$\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{h}^{(k)}_{\mathrm{d}}$] represents the gain of the direct channel, and $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{f}_{\mathrm{k}}]$ is the cross product of the direct channel and channel through the subsurface designed for UE $k$. Higher element spacing at the BS increases the SNR, due to the $\mathbf{R}_{\mathrm{d}}\mathbf{a}_{\mathrm{b}}||$ component of (19). This term reduces if BS elements are placed closely together, as in [18].
@@
Remark 2
$\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{g}_{\mathrm{k}}]$ represents the cross products of the direct channel and the channels through subsurfaces designed for other UEs. The complexity of the summation term in (20) makes it difficult to interpret in detail, but the size of the term is increased, as expected, by alignment of the direct and RIS paths. Note that the summation of a large number of phases in $\mathbf{a}^{(k)}_{\mathrm{r},s}$ and $\mathbf{a}^{(k)}_{\mathrm{ur},s}$ is likely to cause significant cancellation, so this is not a dominant term.
@@
$\times\!{}_{1}F_{1}\!\left(\!\frac{1}{2},2,-\frac{|\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{a}_{\mathrm{d}}^{(k)}|^{2}\kappa_{\mathrm{d}}^{(k)}}{\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{R}_{\mathrm{d}}^{(k)}\mathbf{a}_{\mathrm{b}}^{(k)}}\!\right)\!\sum_{s\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}\bigg[\!\Big(\mathbf{a}^{(k)}_{\mathrm{ur},s,j}\!-\!\rho^{(k)}_{ks}$
$\times\mathbf{a}_{\mathrm{ur},k,i}^{(k)}\Big){}_{1}F_{1}\!\!\left(\!\frac{3}{2},1,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!+\!\frac{3}{2}\rho^{(k)}_{ks}{}_{1}F_{1}\!\!\left(\!\frac{5}{2},2,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}}\!\bigg]$
$\times\!\left(\!\kappa_{\mathrm{rb}}^{(k)}\mathbf{a}^{(k)}_{\mathrm{r},k,i}\mathbf{a}^{(k)*}_{\mathrm{r},s,j}\!+\!r_{ks}^{(k)}\right)\!(\mathbf{C}(s))_{j,j}\,\mathrm{e}^{-j\angle\mathbf{a}_{\mathrm{r},k,i}^{(k)}},$ (22)
@@
Remark 3
$\mathbb{E}[\mathbf{f}_{\mathrm{k}}^{\dagger}\mathbf{g}_{\mathrm{k}}]$ represents the cross products of the channel through the subsurface designed for the UE and the channels through subsurfaces designed for other UEs. As in Remark 2, the main feature of (22) is that the summation of large numbers of weighted phases tends to cause considerable cancellation, reducing the importance of this term.
@@
$\times\!\cos(n\phi_{f})\frac{\epsilon_{n}|\rho^{(k)}_{kk}|^{2m-n}}{m!(m-n)!(n!)^{2}}\!\left(\!\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}-2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\!\right)^{\!\!n}\!\!\$
$\times\!\Gamma^{2}\!\!\left(\!\!m\!+\!\frac{3}{2}\!\right)\!{}_{1}F_{1}^{2}\!\!\left(\!m\!+\!\frac{3}{2},n\!+\!1,\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}-2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\right)\!,$ (25)
@@
Remark 4
$\mathbb{E}[\mathbf{f}_{\mathrm{k}}^{\dagger}\mathbf{f}_{\mathrm{k}}]$ represents the gain of the channel through the subsurface designed for UE $k$. It includes Gaussian hypergeometric functions, which increase monotonically with the final argument. The final argument increases with the spatial correlation between subsurface elements, and this increase occurs when subsurface elements are placed closer together. Hence, placing the subsurface elements close together actually increases the mean SNR.
@@
$\quad\times\mathrm{e}^{-j\angle\mathbf{a}_{\mathrm{r},s,i}^{(s)}}\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{r},s,j}^{(s)}}\Big(\kappa_{\mathrm{ur}}^{(k)}\mathbf{a}_{\mathrm{ur},s,i}^{(k)*}\mathbf{a}_{\mathrm{ur},s,j}^{(k)}+\rho^{(k)}_{ss}\Big)\Big(\kappa_{\mathrm{rb}}^{(k)}$
$\quad\times\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},s,j}^{(k)*}+r^{(k)}_{ss}\Big)\!+\!\!\sum_{s\neq t\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}(\mathbf{C}^{*}\!(s)\!)_{i,i}(\mathbf{C}(t)\!)_{j,j}$
$\quad\times\!\Big(\!\kappa_{\mathrm{ur}}^{(k)}\mathbf{a}_{\mathrm{ur},s,i}^{(k)*}\mathbf{a}_{\mathrm{ur},t,j}^{(k)}\!+\!\rho^{(k)}_{st}\!\Big)\!\Big(\!\kappa_{\mathrm{rb}}^{(k)}\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},t,j}^{(k)*}\!+\!r^{(k)}_{st}\!\Big)\!\Bigg$]\!, (26)
@@
$\times\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}\frac{\epsilon_{n}\kappa_{0}^{2m+n}}{2^{m+n}m!\Gamma(m\!+\!n\!+\!1)}\left(\mathcal{F}(3)\mathrm{e}^{jn\phi}\!+\!\mathcal{F}(1)\mathrm{e}^{-jn\phi}\right)\!,$ (27)
@@
$\!\!\!\times\!\!{}_{1}F_{1}^{2}\!\!\left(\!\!n\!+\!m\!+\!\frac{x}{2},\!n\!-\!1\!+\!x,\!\frac{\zeta^{2}\!\kappa_{\mathrm{ur}}^{(s)}\!\!\left(\!1\!+\!|\rho_{ss}^{(s)}|^{2}\!-\!2\mu_{c}\!\right)\!\!\!\left(\!1\!-\!\rho_{ss}^{(s)2\!}\right)}{4\left(1\!+\!\kappa_{\mathrm{ur}}^{(s)}\right)}\!\!\right)\!\!,$ (28)
@@
Remark 5
$\mathbb{E}[\mathbf{g}_{\mathrm{k}}^{\dagger}\mathbf{g}_{\mathrm{k}}]$ represents the sum of the gains of the channels through subsurfaces designed for other UEs. Increased correlation between elements of any subsurface is still beneficial to UE $k$, again due to the presence of Gaussian hypergeometric functions that increase monotonically with the spatial correlation of RIS elements. Therefore, it is beneficial to all UEs if all RIS elements are closely collocated.
Remark 6
As the terms in (23) and (26) are quadratic, and thus typically larger than the cross product terms, a change in spatial correlation at the RIS is more influential to the overall SNR than a change in spatial correlation at the BS.
Combining (18), (19), (20), (22), (23) and (26) with (17) gives the complete mean SNR.
@@
[…]
@@
Remark 7
The contribution in (30) is enhanced by alignment between the scattering in the direct link and LoS component of the RIS link. Specifically, alignment between the leading eigenvectors of $\mathbf{R}_{\mathrm{d}}^{(k)}$ and $\mathbf{a}_{\mathrm{b}}^{(k)}$ increases the SNR. In (31), increasing channel power enhances the dominant first term. The second summation term tends to increase with spatial correlation due to the monotonically increasing Gaussian hypergeometric function, while the summation over $\mathbf{A}_{i,j}$ terms is less significant due to the cancellation inherent in summing over phases. The same pattern occurs in (32) with increasing correlation boosting the positive $i=j$ component and $i\neq j$ terms tending to cancel out.